# RoboPhD: Self-Improving Text-to-SQL Through Autonomous Agent Evolution

### v1

@@ 1 Introduction
Text-to-SQL, the task of translating natural language questions into executable SQL queries, remains a fundamental challenge in natural language processing with significant real-world applications. Recent advances have shown that large language models (LLMs) can achieve impressive performance on benchmarks like Spider (spider2018) and BIRD (bird2023), but these approaches typically require extensive manual prompt engineering, careful model selection, and domain expertise.
The ability of AI systems to conduct AI research autonomously has been identified as a crucial capability milestone with profound implications for the pace of AI progress (steinhardt2022ai; cotra2022transformative). Several researchers have argued that AI systems capable of improving themselves or conducting research represent a potential discontinuity in technological development (bostrom2014superintelligence; yudkowsky2008artificial). Recent work has begun to explore this frontier, with systems demonstrating increasing autonomy in scientific discovery (romera2024mathematical), machine learning research (lu2024aiscientist), and algorithm development (fawzi2022discovering).
While well short of a full self-improving system, our work demonstrates a concrete instantiation of AI agents autonomously conducting the research cycle of hypothesis generation, experimentation, and iterative refinement to improve performance on text-to-SQL generation.
RoboPhD’s architecture evolves AI agents composed of two artifacts—a database analysis tool and SQL generation instructions—using an architecture that separates offline analysis from online inference. This pattern generalizes to other tasks with similar structure.
@@ 2.1 Text-to-SQL Systems
The BIRD benchmark (bird2023) has driven significant advances in Text-to-SQL, with top systems achieving (as of mid-September 2025) 71-77% accuracy through diverse approaches. AskData (askdata2025) achieves 77.14% using GPT-4o with data analysis and automatic metadata extraction. CHASE-SQL (chase2024) reaches 76.02% through a multi-agent framework employing divide-and-conquer query decomposition, chain-of-thought reasoning based on query execution plans, and instance-aware synthetic few-shot demonstration generation, with a fine-tuned binary selection LLM for candidate ranking. XiYan-SQL (xiyan2024) achieves 75.63% using a multi-generator ensemble strategy that combines in-context learning with supervised fine-tuning, a novel M-Schema representation for database structures, and a fine-tuned selection model to choose among diverse candidates. CSC-SQL (csc2025) reaches 73.67% by integrating self-consistency and self-correction, using GRPO-based reinforcement learning to train both SQL generation and merge-revision models that reconcile conflicting candidates. Reasoning-SQL (reasoning2025) achieves 72.78% through reinforcement learning with SQL-tailored partial rewards encompassing execution accuracy, syntax validity, and n-gram matching. OpenSearch-SQL (opensearch2025) achieves 72.28% using a multi-agent framework with consistency alignment and a novel SQL-Like intermediate language which together reduce instruction-following failures and hallucinations. OmniSQL (omnisql2025) reaches similar accuracy, 72.06%, by leveraging SynSQL-2.5M, the first million-scale synthetic text-to-SQL dataset containing 2.5 million samples with chain-of-thought solutions across 16,000 databases. GenaSQL (genasql2025) achieves 72.28% through ‘N-rep consistency’, a cost-efficient approach that generates diverse SQL candidates using multiple schema representations without chain-of-thought reasoning or fine-tuning, reducing per-query costs to \$0.039. Distillery (distillery2024) reaches 71.83% by fitting the full schema in the input context window, using augmentation, selection, and correction techniques. CHESS (chess2024) achieves 71.1% using a four-agent architecture comprising an Information Retriever, Schema Selector, Candidate Generator, and Unit Tester.
@@
Our work uniquely focuses on autonomous discovery of these strategies rather than manual design. While prior systems require human experts to engineer prompts and architectures, RoboPhD discovers optimization techniques autonomously. Note that we have experimented with using all of the above papers (with the exception of the BIRD benchmark paper (bird2023)) as inputs to the evolution agent in the research-driven evolution strategy mentioned in Section 5.2.
@@
Prior work on prompt optimization includes APE (ape2022), OPRO (opro2023), DSPy (dspy2023), TextGrad (yuksekgonul2025optimizing), and ACE (zhang2025agentic). While these systems optimize prompts, our approach evolves complete agent artifacts including database analysis tools and SQL generation instructions. Crucially, our system conducts research autonomously without human intervention in the optimization loop. One key aspect of our approach is allowing the Evolutionary agent access to the errors made from previous candidate agents. This is similar to the idea of text gradients explored in the TextGrad paper (yuksekgonul2025optimizing), where the LLM produces a (metaphorical) gradient to describe how to address a particular error. In our system, the error gradients are available for the agent to choose how to incorporate them into the next candidate agent.
@@
Evolutionary algorithms have been applied to neural architecture search (nas2019) and hyperparameter optimization (loshchilov2016cmaes; jaderberg2017pbt). We extend this to evolving complete AI agents with natural language specifications and tool usage, creating a fully autonomous research system.
@@
The ELO rating system (elo1978), originally developed for chess, has found recent applications in AI evaluation. The Chatbot Arena (zheng2023) uses ELO ratings to create a dynamic leaderboard of LLMs based on human preferences, demonstrating ELO’s effectiveness for model comparison. While these systems use ELO for passive evaluation and ranking, we introduce ELO as an active selection mechanism for evolutionary optimization (see §3.2.5 for our implementation).
@@ 3.1 Conceptual Overview: The Graduate Student Metaphor
RoboPhD operates like a graduate student operating with minimal faculty supervision, focused on reducing error rate. This paper describes an application to the BIRD Text-to-SQL benchmark, but the lack of SQL-specific initial guidance points towards easy transfer to other domains. The evolutionary agent, here implemented as Claude Code [^1] https://claude.com/product/claude-code, wrapping Claude Sonnet 4.5222https://www.anthropic.com/claude-sonnet-4-5-system-card or Opus 4.5333https://www.anthropic.com/claude-opus-4-5-system-card, can be interpreted as a doctoral student attempting to build a complete AI-powered text-to-SQL system, analogous to prior efforts on the BIRD benchmark.
@@ 3.2.3 Online SQL Generation Phase
Agentic Query Result Validation: Text-to-SQL systems face two critical failure modes: syntax errors and semantic errors. Syntax errors prevent execution, while semantic errors produce incorrect results. Prior work typically addresses these challenges through multiple generation attempts. For instance, OpenSearch-SQL (opensearch2025) generates multiple candidate queries and uses self-consistency voting. Our system employs an agentic answer evaluation approach:
@@ 3.2.6 Experimental Protocol: Selection and Sampling
Agent Selection: In a standard evolutionary iteration, (see Appendix D for details on exceptions), RoboPhD tests three agents selected via a priority system: (1) the previous winner, (2) the newly evolved agent, and (3) a random agent selected from the top 2 performers by ELO (other than the winner). This balance ensures thorough testing while maintaining the evolutionary pressure.
@@ 4.1 Experimental Setup
We evaluate on the BIRD benchmark (bird2023), a large-scale cross-domain dataset with 69 training databases and 11 development databases. Each evolution iteration (using only Haiku 4.5 LLMs for SQL generation) costs approximately \$2 in Claude API for the calls and runs for 22 minutes on average, enabling rapid experimentation cycles. Note that evolution is accomplished via Claude Code (using either Sonnet or Opus 4.5 LLMs) and does not incur direct costs, but rather uses up a weekly usage quota. Empirically, each 30-iteration run uses about 12% of the weekly usage quota of a Claude Max plan. [^4] https://www.claude.com/pricing/max Computational requirements are modest. Experiments were conducted on either a MacBook Pro (48GB RAM) or an Azure VM with 8 vCPUs and 32GB RAM.
@@ 5.3 Broader Applications
Other Text-to-SQL benchmarks. While we evaluate on BIRD, the RoboPhD framework is benchmark-agnostic. Spider 2.0 (spider2) presents a natural extension, with its enterprise-scale databases and more complex queries requiring multi-step reasoning. BIRD-Critic (swesql) offers another direction, focusing on debugging and fixing incorrect SQL queries rather than generation from scratch.
@@ 5. Fall back to manual schema extraction using SQLite CLI:
        print(f"Schema extraction complete - wrote {len(schema_statements)} DDL statements")

### v2

@@ 1 Introduction
Text-to-SQL, the task of translating natural language questions into executable SQL queries, remains a fundamental challenge in natural language processing with significant real-world applications. Recent advances have shown that large language models (LLMs) can achieve impressive performance on benchmarks like Spider (Yu et al., 2018) and BIRD (Li et al., 2023), but these approaches typically require extensive manual prompt engineering, careful model selection, and domain expertise.
The ability of AI systems to conduct AI research autonomously has long been identified as a crucial capability milestone, with the potential to create a discontinuity in technological development (Good, 1965; Chalmers, 2010). Recent work has begun to explore this frontier, with systems demonstrating increasing autonomy in scientific discovery (Romera-Paredes et al., 2024), machine learning research (Lu et al., 2024), and algorithm development (Fawzi et al., 2022).
While well short of a full self-improving system, our work demonstrates a concrete instantiation of AI agents autonomously conducting the research cycle of hypothesis generation, experimentation, and iterative refinement to improve performance on text-to-SQL generation. RoboPhD evolves AI agents composed of two artifacts—a database analysis tool and SQL generation instructions—separating offline analysis from online inference, a pattern that generalizes to other tasks with similar structure.
@@ 2.1 Text-to-SQL Systems
The BIRD benchmark (Li et al., 2023) has driven significant advances in Text-to-SQL, with top systems achieving (as of mid-September 2025) 71-77% accuracy through diverse approaches. AskData (Shkapenyuk et al., 2025) achieves 77.14% using GPT-4o with data analysis and automatic metadata extraction. CHASE-SQL (Pourreza et al., 2024) reaches 76.02% through a multi-agent framework employing divide-and-conquer query decomposition, chain-of-thought reasoning based on query execution plans, and instance-aware synthetic few-shot demonstration generation, with a fine-tuned binary selection LLM for candidate ranking. XiYan-SQL (Liu et al., 2025) achieves 75.63% using a multi-generator ensemble strategy that combines in-context learning with supervised fine-tuning, a novel M-Schema representation for database structures, and a fine-tuned selection model to choose among diverse candidates. CSC-SQL (Sheng and Xu, 2025) reaches 73.67% by integrating self-consistency and self-correction, using GRPO-based reinforcement learning to train both SQL generation and merge-revision models that reconcile conflicting candidates. Reasoning-SQL (Pourreza et al., 2025) achieves 72.78% through reinforcement learning with SQL-tailored partial rewards encompassing execution accuracy, syntax validity, and n-gram matching. OpenSearch-SQL (Xie et al., 2025) achieves 72.28% using a multi-agent framework with consistency alignment and a novel SQL-Like intermediate language which together reduce instruction-following failures and hallucinations. OmniSQL (Li et al., 2025a) reaches similar accuracy, 72.06%, by leveraging SynSQL-2.5M, the first million-scale synthetic text-to-SQL dataset containing 2.5 million samples with chain-of-thought solutions across 16,000 databases. GenaSQL (Dönder et al., 2025) achieves 72.28% through ‘N-rep consistency’, a cost-efficient approach that generates diverse SQL candidates using multiple schema representations without chain-of-thought reasoning or fine-tuning, reducing per-query costs to \$0.039. Distillery (Maamari et al., 2024) reaches 71.83% by fitting the full schema in the input context window, using augmentation, selection, and correction techniques. CHESS (Talaei et al., 2024) achieves 71.1% using a four-agent architecture comprising an Information Retriever, Schema Selector, Candidate Generator, and Unit Tester.
@@
Our work uniquely focuses on autonomous discovery of these strategies rather than manual design. While prior systems require human experts to engineer prompts and architectures, RoboPhD discovers optimization techniques autonomously. Note that we have experimented with using all of the above papers (with the exception of the BIRD benchmark paper (Li et al., 2023)) as inputs to the evolution agent in the research-driven evolution strategy mentioned in Section 5.2.
@@
Prior work on prompt optimization includes APE (Zhou et al., 2022), OPRO (Yang et al., 2023), DSPy (Khattab et al., 2023), TextGrad (Yuksekgonul et al., 2025), and ACE (Zhang et al., 2025). While these systems optimize prompts, our approach evolves complete agent artifacts including database analysis tools and SQL generation instructions. Crucially, our system conducts research autonomously without human intervention in the optimization loop. One key aspect of our approach is allowing the Evolutionary agent access to the errors made from previous candidate agents. This is similar to the idea of text gradients explored in the TextGrad paper (Yuksekgonul et al., 2025), where the LLM produces a (metaphorical) gradient to describe how to address a particular error. In our system, the error gradients are available for the agent to choose how to incorporate them into the next candidate agent.
@@
Evolutionary algorithms have been applied to neural architecture search (Pham et al., 2019) and hyperparameter optimization (Loshchilov and Hutter, 2016; Jaderberg et al., 2017). We extend this to evolving complete AI agents with natural language specifications and tool usage, creating a fully autonomous research system.
@@
The ELO rating system (Elo, 1978), originally developed for chess, has found recent applications in AI evaluation. The Chatbot Arena (Zheng et al., 2023) uses ELO ratings to create a dynamic leaderboard of LLMs based on human preferences, demonstrating ELO’s effectiveness for model comparison. While these systems use ELO for passive evaluation and ranking, we introduce ELO as an active selection mechanism for evolutionary optimization (see §3.2.5 for our implementation).
@@ 3.1 Conceptual Overview: The Graduate Student Metaphor
RoboPhD is analogous to a graduate student operating with minimal faculty supervision, focused on reducing error rate. This paper describes an application to the BIRD Text-to-SQL benchmark, but the lack of SQL-specific initial guidance points towards easy transfer to other domains. The evolutionary agent, here implemented as Claude Code [^1] https://claude.com/product/claude-code, wrapping Claude Sonnet 4.5222https://www.anthropic.com/claude-sonnet-4-5-system-card or Opus 4.5333https://www.anthropic.com/claude-opus-4-5-system-card, can be interpreted as a doctoral student attempting to build a complete AI-powered text-to-SQL system, analogous to prior efforts on the BIRD benchmark.
@@ 3.2.3 Online SQL Generation Phase
Agentic Query Result Validation: Text-to-SQL systems face two critical failure modes: syntax errors and semantic errors. Syntax errors prevent execution, while semantic errors produce incorrect results. Prior work typically addresses these challenges through multiple generation attempts. For instance, OpenSearch-SQL (Xie et al., 2025) generates multiple candidate queries and uses self-consistency voting. Our system employs an agentic answer evaluation approach:
@@ 3.2.6 Experimental Protocol: Selection and Sampling
Agent Selection: In a standard evolutionary iteration (see Appendix D for details on exceptions), RoboPhD tests three agents selected via a priority system: (1) the previous winner, (2) the newly evolved agent, and (3) a random agent selected from the top 2 performers by ELO (other than the winner). This balance ensures thorough testing while maintaining the evolutionary pressure.
@@ 4.1 Experimental Setup
We evaluate on the BIRD benchmark (Li et al., 2023), a large-scale cross-domain dataset with 69 training databases and 11 development databases. Each evolution iteration (using only Haiku 4.5 LLMs for SQL generation) costs approximately \$2 in Claude API for the calls and runs for 22 minutes on average, enabling rapid experimentation cycles. Note that evolution is accomplished via Claude Code (using either Sonnet or Opus 4.5 LLMs) and does not incur direct costs, but rather uses up a weekly usage quota. Empirically, each 30-iteration run uses about 12% of the weekly usage quota of a Claude Max plan. [^4] https://www.claude.com/pricing/max Computational requirements are modest. Experiments were conducted on either a MacBook Pro (48GB RAM) or an Azure VM with 8 vCPUs and 32GB RAM.
@@ 5.3 Broader Applications
Other Text-to-SQL benchmarks. While we evaluate on BIRD, the RoboPhD framework is benchmark-agnostic. Spider 2.0 (Lei et al., 2024) presents a natural extension, with its enterprise-scale databases and more complex queries requiring multi-step reasoning. BIRD-Critic (Li et al., 2025b) offers another direction, focusing on debugging and fixing incorrect SQL queries rather than generation from scratch.
@@ 5. Fall back to manual schema extraction using SQLite CLI:
        print(f"Wrote {len(schema_statements)} DDL statements")